# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v5

@@ Abstract
This paper achieves noteworthy progress in the realm of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo challenges. Initially, we introduce Lico-Net, a novel baseline model that resolves RPM problems with remarkable accuracy. Leveraging this foundation, we advance with the D3C approach, which advocates representing the underlying concepts in abstract reasoning problems through distributions. This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks. To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined yet precise solution. Furthermore, we propose the D2C method, redefining conceptual boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts. Finally, we extend our methodology to D4C, employing adversarial techniques to refine conceptual boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges. Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
@@
Abstract reasoning, RPM problem, Bongard-logo problem.
@@ IV-A2 Reasoning module of Lico-Net
Where $b$ represents the batchsize for training the Lico-Net, and $a$ denotes the count of minimal reasoning units within a single RPM case, which amounts to three in RAVEN and six in PGM. $I$ represents the identify matrix. And $d$ represents the dimension of unit vectors $U_{ij}$. The introduction of the covariance matrix as an additional loss function is an effective technique inherited from Triple-CFN, embodying its essence. The entire feedforward process can be depicted in figure 8.
@@ IV-B The baseline for Bongard-Logo problem
The cases designers of Bongard-logo drew from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. This classification allows us to abstractly represent the distribution of the primary group in each Bongard-logo case as $p_{i}(x|\,y)$, while the distribution of the auxiliary group is denoted as $q_{i}(x|\,y)$. Here, $y$ denotes the reasoning type of the case, which can take values from the set FF, BA, HD, and $i$ represents the case index, ranging from $1$ to $n$, where $n$ is the total number of cases. Additionally, we denote the Bongard-Logo images as $x_{ij}$. Specifically, $\{x_{ij}|\,j\in[1,6]\}$ represents images in the $i$-th primary group, while $\{x_{ij}|\,j\in[8,13]\}$ represents images in the $i$-th auxiliary group. Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group.
We copy a baseline from Triple-CFN, which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. The InfoNCE loss function[41] can be mathematically expressed as:
@@
Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group. $m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$ in this paper.
Through the strategic employment of the InfoNCE function, we can circumvent the need to explicitly estimate the distributions $p_{i}^{\prime}(z|\,y)$ and $q_{i}^{\prime}(z|\,y)$. Concurrently, this approach encourages enhanced similarity among the representations of the primary group encoded by the network $f_{\theta}(z|\,x)$ and guarantees a greater degree of mutual exclusion between the representations of the primary and auxiliary groups. This process aligns with the underlying logic of Bongard-logo. InfoNCE can be construed as a loss function devised for the clustering of vector groups.
This paper reproduces the experiments for the baseline in Triple-CFN conducted based model ResNet50, which are individually implemented on the Bongard-logo database based on ResNet18. It focuses on the three concepts: ‘FF’, ‘BA’, and ‘HD’ for training, while concentrate on four test concept: ‘FF’, ‘BA’, ‘NV’ and ‘CM’. Subsequently, a combined database was created by merging these three individual databases, and experiments were performed on this combined database. The experimental results are presented in Table I. The feedforward process is shown in figure 9.
The findings indicate that when presented with a combination of the three databases, the convolutional deep model experiences confusion among the three distinct human concepts. The varying degrees of reduction in the model’s reasoning accuracy on the FF and BA problems, as well as the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion is not attributable to insufficient network capacity, as even the ResNet50 employed in this study, which is more capacious than the ResNet18 utilized in Triple-CFN’ s baseline, falls victim to the same predicament.
@@
ResNet50 on Bongard-logo 88.9 98.1 76.1 76.0
@@
This paper posits that, in the Bongard-Logo problem, the determination of image attributes relies heavily on the primary and auxiliary group samples within each case, which poses a fundamental challenge. It is precisely this multifaceted attribution of image properties that confuses deep networks when dealing with the three high-dimensional concepts inherent in the Bongard-Logo problem. Conversely, the RPM problem exhibits consistency in its recognition of image attributes. For instance, a white medium-sized triangle retains its attribute identification as such, regardless of its appearance in any RPM case option pool.
@@ V-A D3C for Bongard-Logo
The ideal approach to address the Bongard-Logo problem involves training an encoder capable of encoding images from a Bongard-Logo case into representations, denoted as $z$. In this context, the representations of positive instances (samples in primary group) belong to the distribution $p(z)$, while the representations of negative instances (samples in primary group) belong to $q(z)$. Importantly, there should be minimal overlap between $p(z)$ and $q(z)$ in terms of cross-measure. Therefore, this paper proposes the D3C method to enhance the performance of encoders in the form of $f(z|x)$ on Bongard-Logo problems. Notably, this brings improvements for models such as Triple-CFN and ResNet50.
Specifically, this paper hypothesizes that the representation of the primary group samples in a Bongard-Logo case follows a Gaussian distribution, while the representation of the auxiliary group samples follows a mixture of Gaussian distributions. We utilize an arbitrary backbone network $f(z|\,x)$ to approximate the distribution of image representations within Bongard-Logo cases. To achieve this, we modify the network to output parameters $\mu$ and $\sigma$ instead of a single vector, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The encoded vector is then divided into two equal-length segments, which we interpret as the mean and the logarithm of the variance, respectively. This approach normalizes the output vectors into Gaussian distributions and facilitates the distribution of image representations in Bongard-Logo cases. Our method bears resemblance to the Variational Autoencoder (VAE) framework. We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
With the estimated distributions of representations for each image in the Bongard-Logo cases, we compute the Sinkhorn distance between these distributions using reparameterization technique. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 10. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
@@
Where $b$ is the batchsize of training. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
@@ V-B D3C on Lico-Net for RPM
Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule. Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
@@ VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is significantly influenced by specific contextual concepts associated with each individual case. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem.
To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts. It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. In other words, each case is uniquely categorized based on its distinct set of concepts. By establishing a one-to-one correspondence between image styles and these redefined concepts, we aim to enable the network to seamlessly infer concepts from the pixel arrangements of images. This, in turn, aims to mitigate the impact of multiple attribute identifications on the solution of the Bongard-Logo problem.
To foster concept antagonism between different problem domains, we have revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture. This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse problem sets.
@@ VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
By augmenting the traditional sampling methods used for mini-batch training with an additional ancestral sampling procedure, this study identifies mutually exclusive pairs of Bongard-logo problems for training purposes. This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
@@
In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative instances in the current case. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
@@ VI-B D2C on RPM probelm
In summary, D2C is a straightforward approach. In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases. Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM. It is worth emphasizing that the D2C and D3C methodologies are not mutually exclusive. Specifically, D2C endeavors to redefine concept boundaries for reasoning problems. According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.

### v9

@@ Abstract
This paper achieves noteworthy progress in the realm of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo challenges. Initially, we introduce Lico-Net, a novel baseline model that resolves RPM problems with remarkable accuracy. Leveraging this foundation, we advance with the D3C approach, which advocates representing the underlying concepts in abstract reasoning problems through distributions. This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks. To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined solution. Furthermore, we propose the D2C method, redefining concept boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts. Finally, we extend our methodology to D4C, employing adversarial techniques to refine concept boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges. Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
@@
Abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Adversarial Learning.
@@ IV-A2 Reasoning module of Lico-Net
Where $b$ represents the batchsize for training the Lico-Net, and $a$ denotes the count of minimal reasoning units within a single RPM case, which amounts to three in RAVEN and six in PGM. $I$ represents the identify matrix. And $d$ represents the dimension of unit vectors $U_{ij}$. The entire feedforward process of Lico-Net can be depicted in figure 8.
@@ IV-B The baseline for Bongard-Logo problem
The cases designers of Bongard-logo drew from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. It is worth noting that the concept HD is decomposed into NV and CM during testing. This classification allows us to abstractly represent the distribution of the primary group in each Bongard-logo case as $p_{i}(x|\,y)$, while the distribution of the auxiliary group is denoted as $q_{i}(x|\,y)$. Here, $y$ denotes the reasoning type of the case, which can take values from the set FF, BA, HD, and $i$ represents the case index, ranging from $1$ to $n$, where $n$ is the total number of cases. Additionally, we denote the Bongard-Logo images as $x_{ij}$. Specifically, $\{x_{ij}|\,j\in[1,6]\}$ represents images in the $i$-th primary group, while $\{x_{ij}|\,j\in[8,13]\}$ represents images in the $i$-th auxiliary group. Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group.
We copy a baseline from Triple-CFN [32], which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. The InfoNCE loss function[41] can be mathematically expressed as:
@@
Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the representations in primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group. $m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$.
This approach fosters a higher degree of similarity between $z_{pri}$ and $\tilde{z}_{pri}$ compared to the similarity between $z_{pri}$ and $z_{au{x_{m}}}$, thereby emphasizing the distinction between primary and auxiliary representations. This process aligns with the underlying logic of Bongard-logo. InfoNCE can be construed as a loss function devised for the clustering of vector groups.
This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model. The baseline focuses on learning four concepts in Bongard-Logo — ’FF’, ’BA’, ’NV’, and ’CM’ — separately or in combination [32]. To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments. The experimental results are presented in Table I, while the feedforward process is illustrated in Figure 9.
The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts. The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, coupled with the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
@@
ResNet50 on Bongard-logo 88.9 98.0 76.1 76.0
@@
After conducting backbone substitution experiments, this paper aligns with the perspective of Triple-CFN [32]. Triple-CFN argues that the identification of image attributes in the Bongard-Logo problem relies heavily on primary and auxiliary samples in specific cases [32], posing challenges due to this context-dependent attribute determination. In other words, the reasoning identity of these images is contingent upon context-specific concepts, leading to multifaceted interpretations of image attributes. Consequently, deep networks encounter confusion when processing high-dimensional Bongard-Logo concepts [32]. In contrast, the RPM problem exhibits consistency in its recognition of image attributes, such that a white medium-sized triangle retains its identity regardless of its appearance in any RPM case option pool.
@@ V-A D3C for Bongard-Logo
D3C proposes that distributions are more suitable as abstract representations of concepts compared to vectors. Therefore, D3C decides to estimate the concept in each image of a Bongard-Logo case using a deep network in the form of $f(z|\,x)$ and represent the concept with a distribution. Finally, the distances between these distributions are measured and used as the distances between concepts. These distances can be naturally applied to solve the Bongard-Logo problem. D3C is a plug-and-play method targeted at solving Bongard-Logo and similar problems for any network in the form of $f(z|\,x)$. Notably, this brings improvements for models such as Triple-CFN and ResNet50.
In practice, this paper hypothesizes that the concepts of the image samples in a Bongard-Logo case follows a Gaussian distribution. We utilize an arbitrary backbone network $f(z|\,x)$ to excract the concepts of images within Bongard-Logo cases in form of approximating distribution. To achieve this, we modify the network to output two vectors $\mu$ and $\sigma$ instead of a single vector $z$, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The encoded vector $z$ is then divided into two equal-length segments $\mu$ and $\log(\sigma^{2})$, which we interpret as the mean and the logarithm of the variance, respectively. This approach normalizes the output vectors into Gaussian distributions and represents concepts as distributions in Bongard-Logo cases. Our method bears resemblance to the Variational Autoencoder (VAE) framework[8]. We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
With the estimated concepts for each image in the Bongard-Logo cases as distributions, we compute the Sinkhorn distance between these distributions using reparameterization technique[8]. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 10. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
@@
Where $b$ is the batchsize of training. The $\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\}$ represents $S$ samples drawn from the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ using the reparameterization trick, where the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ is obtained by mapping the Bongard-Logo image $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$ through the network $f(\mu,\sigma|\,x)$. The network $f(\mu,\sigma|x)$ is derived from the network $f(z|x)$ by D3C Method. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
@@ V-B D3C on Lico-Net for RPM
Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule. Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns. This approach is precisely the original idea behind D3C, which posits that describing concepts using distributions is more appropriate than using vectors. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
@@ VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32]. Distinct concepts require the network to encode specific pixel configurations into differentiated attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32]. This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32]. In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem. We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts. It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. In other words, each case is categorized and solved based on its distinct concepts. By establishing a one-to-one correspondence between problem cases and these redefined concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case. This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
To foster concept antagonism between different Bongard-Logo cases, we have revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture. This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse Bongard-Logo cases.
@@ VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
By incorporating an additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes. This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
@@
In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative samples in the current case. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
@@ VI-B D2C on RPM probelm
In summary, D2C is a straightforward approach. In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases. Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM. In a nutshell, D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases. It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection. Finally, the D2C and D3C methodologies are not mutually exclusive. Specifically, D2C endeavors to redefine concept boundaries for reasoning problems. According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.